OpenStudy (anonymous):

Working together, paul and daniel can pick forty bushels of apples in 4.95 hours.had he done it alone it would have taken daniel 9 hours. Find how long it would take paul to do it alone?

OpenStudy (anonymous):

\[r_1+r_2=\frac{1}{4.95}\] \[r_1=\frac{1}{9}\] solve for \[r_2\]

OpenStudy (anonymous):

actually the question is "how long?" and i solved for the rate. you are looking for time which is \[\frac{1}{r_2}\]

OpenStudy (anonymous):

reasoning as follows: their combined rate is 4.95 meaning they do \[\frac{1}{4.95}\] of the job per hour. by the same reasoning daniel's rate is \[\frac{1}{9}\] meaning he does one-ninth of the job per hour.

OpenStudy (anonymous):

so you can write \[\frac{1}{9}+r_2=4.95\] and solve for the second rate \[r_2\] or you can write \[\frac{1}{9}+\frac{1}{x}=4.95\] and x will be your answer

OpenStudy (anonymous):

excuse me the right side of these should be \[\frac{1}{4.95}\] not 4.95

OpenStudy (anonymous):

last one is \[\frac{1}{9}+\frac{1}{x}=\frac{1}{4.95}\]

OpenStudy (anonymous):

i think it is 11 when you solve. you can check. answer is 11 hours